OpenStudy (anonymous):

what is equivalent to -4.5?

OpenStudy (anonymous):

What is the context? -9/2?

OpenStudy (anonymous):

=-45/10 = -9/2

OpenStudy (anonymous):

-4.5 is equivalent to \[-4\frac{1}{2} \]

OpenStudy (anonymous):

also -9/2